An ideal gas expands isothermally, performing 3.40x103 J of work in the process.

Calculate (a) the change in the internal energy of the gas, and (b) the heat

absorbed during the expansion.


1
Expert's answer
2022-01-20T10:07:43-0500

(a) Since the temperature is constant and the change in internal energy depends on the change in temperature, "\Delta U=0."

(b) According to the first law of thermodynamics, "Q=\Delta U+W=0+3.4·10^3=3.4·10^3\text{ J}."